OpenStudy (anonymous):

same as before... saifoo? slaaibak? help mee? The sum of the digits of a two digit number is one third of the number. The units digit is 5 more than the 10s digit. What is the number?

OpenStudy (anonymous):

I think its u+5=t t+u=10t+u/3 but im not really sure.

OpenStudy (anonymous):

HI MY BUDDY SAIFOO!!!!

OpenStudy (saifoo.khan):

Hello Cloverpetal... =

OpenStudy (slaaibak):

10t + u = 3(t + u) t = u - 5 10(u + 5) + u = 3(u + 5 + u) 10u - 50 + u = 6u + 15 5u = 35 u = 7 t = 2 27.

OpenStudy (anonymous):

moneybird thanks but thats not really what I was looking for.

OpenStudy (anonymous):

slaaibak, why did you multiply that second part of the equation by 3?

OpenStudy (slaaibak):

Sorry, there are typos in my equations, i'll fix now

OpenStudy (anonymous):

thanks

OpenStudy (slaaibak):

10t + u = 3(t + u) t = u - 5 10(u - 5) + u = 3(u - 5 + u) 10u - 50 + u = 6u - 15 5u = 35 u = 7 t = 2 27 is the number.

OpenStudy (slaaibak):

I multiplied it by three, since the sum of the digits is 3 times smaller than the actual number, so by multiplying it by 3, it's equal to the number

OpenStudy (anonymous):

why did you subtract five from u when u is five more than the other one?

OpenStudy (slaaibak):

if you are 15 and you are 5 years older than me, my age = 15 - 5

OpenStudy (anonymous):

while you are at it, can you help me with a few more?

OpenStudy (slaaibak):

I'll do one more, have to go sleep after this

OpenStudy (anonymous):

okay... THe sum of three digits is nine. The tens digit is 1 more than the hundreds digit. When the digits are reversed, the new number is 99 less than the original number. What is the original number?

OpenStudy (slaaibak):

x + y + z = 9 y = x + 1 100z + 10y + x + 99 = 100x + 10y + z 99z - 99x = -99 z - x = -1 Now we have: z - x = -1 x + y + z = 9 y = x + 1 In x y z c form: x + 0y - z + 1 = 0 x + y + z - 9 = 0 x - y + 0z + 1 = 0 x=3 y=4 z=2 342 For some reason I got negative numbers, but it works when I make them positive. I'm way too tired to look into that now. x=3 y=4 z=2

OpenStudy (slaaibak):

Should be: x + y + z = 9 x - y = -1 x - z = 1 Solving gives x=3, y=4, z=2